OpenStudy (anonymous):

An airplane takes 1 hour longer to go a distance of 816 miles flying against a headwind than on the return trip with a tailwind. If the speed of the wind is 50 miles per hour, find the speed of the plane in still air.

OpenStudy (amistre64):

you can use distance equations to help solve this

OpenStudy (amistre64):

distance = speed x time the distance remains constant, the speeds remain constant, its the time and the interaction that vary (a+w) t = d (a-w) (t+1) = d

OpenStudy (anonymous):

a=? w=?

OpenStudy (amistre64):

w is wind speed, a is airplane speed, at least in my head

OpenStudy (amistre64):

now using these 2 equations ... we can work up something that with any luck will give us a relationship that we can solve (a+w) t = d (a-w) (t+1) = d -at -wt = -d at +a -wt -w = d a = w(2t+1) but t = d/(a+w) so .. a = w(2d/(a+w)+1) a = 2wd/(a+w) +w a(a+w) = 2wd +w(a+w) a(a+w)-w(a+w) = 2wd a^2 +aw -aw -w^2 = 2wd a^2 -w^2 = 2wd a^2 = 2wd + w^2

OpenStudy (amistre64):

im sure theres a simpler process .... but i can never recall it off hand

OpenStudy (anonymous):

i need to get the a? isn't?

OpenStudy (amistre64):

yes, our goal is to get the speed of the airplane (a)

OpenStudy (anonymous):

it's not factorable using factoring.. can i use quadratic formula? but the answer may have square roots

OpenStudy (amistre64):

i cant readily see a way to remove the time variable ... i keep coming up with solutions that are not a single point.

OpenStudy (amistre64):

a = 50(2t + 1)

OpenStudy (amistre64):

lets see how well this plays tho, lets assume it takes t=1 to get there a = 50(3) = 150 150+50 = 200, not 816 so my thought is evidently not correct

OpenStudy (anonymous):

aww :(

OpenStudy (amistre64):

maybe this makes better sense?